OpenStudy (anonymous):

what are the x intercepts of the quadratic equation 2x^2-8x+3

OpenStudy (anonymous):

do you know the quadratic formula?

OpenStudy (anonymous):

no i forgot it

OpenStudy (anonymous):

\(\Large x\; =\; \frac {-b\; \pm\; \sqrt {b^2\; -\; 4ac}}{2a} \) a=2 b=-8 c=3

OpenStudy (anonymous):

and this will help me find both x intercepts ?

OpenStudy (anonymous):

yes! to find solutions/zeros/roots/x-int for a quadratic equation: \(a x^2\; +\; b x\; +\; c\; =\; 0 \), we always use the quadratic formula.

OpenStudy (anonymous):

im stuck i at this section i have: \[-(-8)\pm \sqrt{\frac{ 40 }{ 4 }}\] what do i do after this ?

OpenStudy (anonymous):

\(\Large x\; =\; \frac {-(-8)\; \pm\; \sqrt {(-8)^2\; -\; 4(2)(3)}}{2(2)} \\\Large x\; =\; \frac {-(-8)\; \pm\; \sqrt {64\; -\;24}}{4}\\\Large x\; =\; \frac {8\; \pm\; \sqrt {40}}{4}\) break it down \(\Large x\; =\; \frac {8\; +\; \sqrt {40}}{4}\) and \(\Large x\; =\; \frac {8\; -\; \sqrt {40}}{4}\) then solve it, you've got two x-int (ofc since it is a quadratic function)

OpenStudy (anonymous):

since 40 doesn't have a square root how do i solve it ?

OpenStudy (anonymous):

i usually don't ask for the answer but i didn't go to the last class where we discussed this so i am completely lost so could you help me out and just explain your answer so i understand how you got there

OpenStudy (anonymous):

\(\sqrt{40}=2 \sqrt{10}=6.32\) \(\Large {x\; =\; \frac {8\; +\; \sqrt {40}}{4}\ and \ \Large x\; =\; \frac {8\; -\; \sqrt {40}}{4}\\x= \; \frac {4\; +\; \sqrt {10}}{2}=3.58 \ or \ x= \; \frac {4\; -\; \sqrt {10}}{2}=0.42}\) there you go , done :) it's 3.58 or 0.42

OpenStudy (anonymous):

you can use either the decimal or the simplified form of the radical term

OpenStudy (anonymous):

do you have any confusions about this topic?

OpenStudy (anonymous):

no you were a big help thank you so much i really appreciate the help

OpenStudy (anonymous):

can i ask you one more question ?

OpenStudy (anonymous):

sorry i was afk, okay sure, i'll try :)